OpenStudy (anonymous):

a paddle boat can move at a speed of 10km/h in still water. the boat is paddled 12km downstream in a river in the same time it takes to go 6km upstream. what is the speed of the river??

OpenStudy (anonymous):

10 is the speed of the boat c is the current 1/2(10+c) = 10 - c 5 + c/2 = 10 - c 3/2 * c = 5 c = 10/3 let me know what you get, i did this pretty quick

OpenStudy (anonymous):

thats that i got too, thank you!